OpenStudy (anonymous):

A box is to be mailed. The volume in cubic inches of the box can be expressed as the product of its three dimensions: V(x) = x^3 - 16x^2 + 79x - 120. The length is (x - 8). Find the linear expressions for the width and the height. Assume that the width is greater than the height.

OpenStudy (dumbcow):

Divide V(x) by (x-8). easiest to use synthetic division 8 | 1 -16 79 -120 8 -64 120 -------------------- 1 -8 15 0 leaving x^2 -8x +15 Factor, -5*-3=15 and -5-3=-8 --> (x-5)(x-3)

OpenStudy (anonymous):

Thank you. Which one of those would be the width and which would be the height?

OpenStudy (dumbcow):

oh right well width > height which is greater, x-5 or x-3 ?

OpenStudy (anonymous):

oh gosh...umm (x-5)

OpenStudy (dumbcow):

say x=10 x-8 = 2 x-5 = 5 x-3 = 7

OpenStudy (anonymous):

so (x-3) . Thanks a ton!